OpenStudy (anonymous):

Is the sum for the first five terms in a sequence where a_1=3 and r=1/3 1.48? I did the process of 3*1/3=1 etc...

ganeshie8 (ganeshie8):

sum = a(1-r^n)/(1-r) a = 3 r = 1/3 n = 5

OpenStudy (anonymous):

Ah so 4.48. Only 3 away from the right answer.. ha

ganeshie8 (ganeshie8):

ahh yes... u forgot to add the first term 3 i see :)

OpenStudy (anonymous):

Damnit. So close.

ganeshie8 (ganeshie8):

happens ;)

OpenStudy (anonymous):

everyday yo, everyday